The longest side of a triangle is 5 inches longer than the shortest side. The medium side is 4 inches longer than the shortest side. If the perimeter of the triangle is 27 ​inches, what are the lengths of the three​ sides?
Answer by Cromlix(4381)

Hi there,
Shortest side = 'x' m
Longest side = x + 5
Medium side = x + 4
Perimeter = three sides added together
27 ins. = x + (x + 5) + (x + 4)
27 ins. = x + x + 5 + x + 4
27 ins. = 3x + 9
27 - 9 = 3x
3x = 18 ins.
x = 6 ins.
Shortest side = 6 ins.
Longest side = 11 ins.
Medium side = 10 ins.
Hope this helps :-)